OpenStudy (anonymous):

a stone is released from an elevator going up with acceleration ‘ at ’. the acceleration of the stone after the release is

OpenStudy (anonymous):

Neglecting air resistance, the stone should be at the same acceleration at the instant upon releasing the stone out of one's hands.

OpenStudy (anonymous):

The stone is in now in free fall - the only force on it is its weight. So it falls with an acceleration g... From the outside, it is clear that the elevator has an upwards acceleration of \[at m/s ^{2}\] and the stone has a downwards acceleration of 9.8m/s^2 (g=-9.8m/s^2). But from inside of the elevator, it appear that the stone is accelerating more than the usual free fall acceleration. That's because as the stone accelerates downwards, the lift is accelerating upwards towards the stone. The acceleration appears to be g-a = -9.8-at.